OpenStudy (templeguy):

what was the purpose of the arrow scene in shrek? It didn't progress the story or anything.

OpenStudy (matlee):

use the formula \[\prod_{\theta \alpha \le \iota \kappa}^{?}\int\limits_{\sqrt[3]{72}}^{shrek}\]